### v2

# PROFIT: A Specialized Optimizer for Deep Fine Tuning

Abstract
Fine-tuning pre-trained models has become invaluable in computer vision and robotics. Recent fine-tuning approaches focus on improving efficiency rather than accuracy by using a mixture of smaller learning rates or frozen backbones. To return the spotlight to model accuracy, we present PROFIT (Proximally Restricted Optimizer For Iterative Training), one of the first optimizers specifically designed for incrementally fine-tuning converged models on new tasks or datasets. Unlike traditional optimizers such as SGD or Adam, which make minimal assumptions due to random initialization, PROFIT leverages the structure of a converged model to regularize the optimization process, leading to improved results. By employing a simple temporal gradient orthogonalization process, PROFIT outperforms traditional fine-tuning methods across various tasks: image classification, representation learning, and large-scale motion prediction. Moreover, PROFIT is encapsulated within the optimizer logic, making it easily integrated into any training pipeline with minimal engineering effort. A new class of fine-tuning optimizers like PROFIT can drive advancements as fine-tuning and incremental training become increasingly prevalent, reducing reliance on costly model training from scratch.
1 Introduction
Fine-tuning pre-trained models has become widely adopted in solving computer vision and robotics problems. As datasets and models both increase in size, having to train a new model for every new application and setting quickly becomes intractable. Imagine, for example, the cost of having to train a new model every time an autonomous vehicle needs to operate in a new city or every time a camera application needs to recognize a new type of object. The shift towards fine-tuning within the deep learning community has also been accelerated by developments in large foundational models that were trained on vast quantities of data, such as CLIP [40] and DINO [5]. Indeed, hints abound that deep learning is steadily inching towards a new paradigm where only simple models are trained from scratch.
At the same time, fine-tuning a model comes with its own set of challenges. For one, it is known that models tend to readily forget old information when fine-tuned on new information, in a process known as Catastrophic Forgetting [16]. Various mitigation methods have been proposed [30], but require additional data engineering and modifications of the model architecture. The common practice within fine-tuning still largely relies on training on new tasks/data with a smaller learning rate or with a frozen backbone (or both). Parameter-efficient fine-tuning methods such as LoRA [18] introduce learnable adapter modules for transfer learning, but serves mainly to improve fine-tuning efficiency rather than accuracy [2]. We aim to specifically improve the accuracy of fine-tuning in the present work, while keeping the process efficient by not introducing any additional parameters.
One ubiquitous element within deep learning training that is also almost always implemented in a modular fashion is the optimizer. The majority of models set the optimizer to some noncontroversial choice (such as Adam  [24], AdamW [34] or Momentum [47]). Yet, all current optimizers are designed with training-from-scratch in mind, which forces their design to assume minimal structure about the problem setting. In contrast, the fine-tuning setting usually starts from a well-trained, well-converged model that we already trust. So we ask: how do we design an optimizer especially for the case where we start from a converged model from a similar domain? Given the ubiquity and modularity of optimizers within training pipelines, such an optimizer would be immediately applicable and easy to implement within any fine-tuning setting.
Works such as Learning Without Forgetting (LWF) [30] have proposed mitigating catastrophic forgetting by enforcing proximity to an old state of the model, but requires additional data pipelining and model snapshots that serve as additional supervision to keep the model anchored to its old “good” state. Instead of anchoring a model across tasks, we propose a different but equally valid anchoring across time. For each iteration of the optimizer, we can take a step away from equilibrium and balance further steps away from equilibrium with the model’s desire to return to equilibrium using standard methods from the gradient-based multitask learning literature. The result is a system that mimics data-driven anchor methods such as [30] without actually needing data processing and replaces the rigid static anchors of other methods with a dynamically updated flexible one.
Figure 1: Schematic of a gradient update for PROFIT (Proximally Restricted Optimizer For Iterative Training). Unlike standard optimizers, which take successive steps through parameter space, PROFIT keeps track of a “good” model state and collects reference information of how the system tries to move away from the “good” state (denoted by the superscript “r” - for “reference” - in the figure). The functionality of the “good” state is then restored by translating and orthogonalizing gradient updates with respect to that state. These processes allow PROFIT to take heavily regularized update steps and produce superior performance when in challenging fine-tuning settings. The visualization in the bottom row is discussed in Section 4.1.
Specifically, a model converged in some state $\mathbf{x}_{0}$ will proceed along the states $\mathbf{x}_{1},\ldots\mathbf{x}_{t}$ with a standard optimizer. A further update would send the system to $\mathbf{x}_{t+1}$. However, because $\mathbf{x}_{0}$ was a “good state,” the model also would benefit by returning to $\mathbf{x}_{0}$. We thus have two potentially conflicting gradient directions ($\mathbf{\Delta}:=\mathbf{x}_{t}\rightarrow\mathbf{x}_{0}$ and $\mathbf{g}:=\mathbf{x}_{t}\rightarrow\mathbf{x}_{t+1}$), which is a classic multitask learning problem. We borrow from [55] and assign $\mathbf{g}\mapsto\mathbf{g}\perp\mathbf{\Delta}$, where $\perp$ is the orthogonalization operator. We restore the model to state $\mathbf{x}_{0}$ (“translate” operation on the top right of Figure 1) and take a step in the orthogonalized direction $\mathbf{g}$.
The operation of PROFIT (Proximally Restricted Optimizer For Iterative Training) relies on compatibility between the fine-tuning and baseline settings, as it aims to keep the model state close to the baseline dynamically. PROFIT excels at adapting a model to a nearby distribution, making this compatibility a crucial aspect of our approach. This excludes some cases often referred to as fine-tuning (e.g., pre-training on ImageNet, fine-tuning on unrelated datasets), we distinguish these pre-training scenarios from practical fine-tuning under compatibility constraints. We refer to this approach as “proximal fine-tuning”, where the fine-tuning dataset is from a similar distribution to pre-training. Although this constraint may seem limiting, “proximal fine-tuning” is prevalent in machine learning applications. For instance, a self-driving car’s trajectory prediction network may need to train a prediction head for a new scooter type using the same sensors.
We name our method PROFIT, which stands for PROximal FIne Tuning, and a schematic of the process is shown in Figure 1. To the best of our knowledge, PROFIT is one of the first optimizers that is specifically designed to work in the fine-tuning setting. PROFIT provides significant performance boosts in those cases, as we will demonstrate.
Our main contributions are as follows:
  • We present PROFIT, an optimizer designed specifically for fine-tuning a good baseline model, and which can be easily dropped into any deep learning training framework.
  • We prove that PROFIT allows us to train as if we still had access to the old data/setting in an unsupervised way.
  • We experimentally show that PROFIT beats standard fine-tuning methods in a large number of settings, from a low-dimensional toy example to image classification to large-scale motion prediction for autonomous driving.
2 Related work
Multi-Task Learning
For detailed background context in multi-task learning (MTL), we refer the reader to [59]. MTL [59] is an optimization problem where we train a model on multiple tasks to take advantage of the structure of shared neural networks, thus improving generalization. One direction is to use gradient descent method to optimize joint multi-task learning problem. Our setting shares some similarities with that. The difference is that our problem is considered as temporal multi-task learning. Ozan [44] formulates the MTL problem as a multi-objective optimization problem and learns the loss weights that change dynamically. GradNorm [10] attempts to normalize gradients to balance learning of multiple tasks. PCGrad [55] suggests that to mitigate gradient direction conflicts, we should project a task’s gradient onto the normal plane of the gradient of any other tasks where a gradient conflict is present. Our proposed technique is similar to PCGrad, but modifies the core designs specifically for the “proximal fine-tuning” setting.
Adapters
An adapter introduces learnable parameters to help transfer learning. The number of trainable parameters in the adapter is much smaller than the model itself, which makes them appealing in practice. Several strategies have been proposed such as visual prompt tuning [21], side adapters [57], bias tuning [4], and residual adapters [41] for efficient transfer learning. LoRA [18] updates the projection matrices in an attention module by decomposing them to a low-rank matrix. In contrast to our method, adapters require additional parameters and may also assume a specific class of model architectures for their success (e.g: transformers).
Optimizer-based Fine-Tuning
Fine-tuning on particular datasets [27, 9] is a common technique in the era of deep learning. Practitioners often use standard optimizers such as Stochastic Gradient Descent (SGD) [3], Adam [24], and AdamW [34]. Recent architectures such as ViTs [12, 5, 40] or ConvNeXts [32] use AdamW for fine-tuning, while it is also common to use SGD for fine-tuning models like ResNets [17, 26] due to the optimizer’s efficiency. However, SGD and AdamW do not assume that we want to stay close to our model’s start state, and thus lead to models that tend to compromise performance on old data when fine-tuning on new data, also known as catastrophic forgetting  [37]. Our method serves as a regularization approach to bridge the gap in existing optimizers to mitigate this forgetting issue.
Continual Learning
To address this issue, various approaches have been developed, such as regularization-based methods [25, 8, 22, 48, 19], where the goal is to keep learned information of the past tasks during continual learning. In particular, Mirzadeh et al. [19] demonstrated a link between continual learning and multi-task learning, and Farajtabar et al. [14] proposed projecting gradients from new tasks onto the subspace of prior task gradients. Learning Without Forgetting (LWF) [30] proposes storing the old model response on new tasks/data and supervises the new model on these responses using a distillation loss to prevent catastrophic forgetting [36]. Our work draws inspiration from LWF’s key idea of storing a pretrained model’s response, but we aim to avoid the storage of old data/checkpoint/statistics. Instead, we use the network’s response at the initial state of each iteration. Two other directions to address catastrophic forgetting are rehearsal-based methods [42, 6, 33, 7, 43] that directly make use of the old data source, and architecture-based methods that minimize intertask interference through new architectures [35, 45, 29, 52, 53]. These methods add substantial infrastructural overhead, while our approach is attractive in its simplicity of implementation. Since our method is an optimizer by design and does not require any additional storage (unlike the continual learning paradigm), we do not benchmark against continual learning methods.
3 Method
3.1 The PROFIT Optimizer
We describe the complete algorithm of PROFIT in Alg 1. Our method is easy-to-understand and modular, which can be encapsulated within the definition of an optimizer step. Next, we provide a discussion of the hyperparameters and the details of the implementation, and conclude with our assumptions and the limitations of PROFIT.
Algorithm 1 PROFIT: A fine-tuning optimizer
1:Converged model $\mathcal{M}(\mathbf{x};\mathbf{\theta})$ with trainable weights $\mathbf{\theta}$ and input $\mathbf{x}$, to be trained on data from similar domain $\mathbf{X}^{\prime}$ with loss $L$.
2:Initialize reference model weights $\mathbf{\theta}_{\text{ref}}$.
3:Initialize batch size $B$, reference steps $n_{\text{ref}}$, training steps $n_{\text{steps}}$, standard optimizer $\mathbf{O}$ with learning rate $\lambda_{\text{main}}$, and reference optimizer $\mathbf{O}^{\text{(ref)}}$ with learning rate $\lambda_{\text{ref}}$. Each optimizer takes as arguments the current weights and a gradient update direction, producing updated weight values.
4:
5:for $n_{\text{step}}$ steps do:
6:     $\mathbf{\theta}_{\text{ref}}\leftarrow\mathbf{\theta}$triangleright Save the model state.
7:     for $n_{\text{ref}}$ steps do
8:         Take new $B$ examples from $\mathbf{X}^{\prime}$ and calculate gradients $\mathbf{g}:=\nabla_{\mathbf{\theta}}L$.
9:         Take one step with reference optimizer $\theta\leftarrow\mathbf{O}^{\text{(ref)}}(\theta,\mathbf{g})$.
10:     end for
11:
12:     Calculate $\Delta=\mathbf{\theta}-\mathbf{\theta}_{\text{ref}}$. $\triangleright$ Calculate total displacement during reference steps.
13:     Find $\mathbf{g}:=\nabla_{\theta}L$ for a new batch as in Line 4.
14:     Calculate dot product $\omega=\langle\Delta,\mathbf{g}\rangle$.
15:     if $\omega<0$ then:
16:         $\mathbf{g}\leftarrow\mathbf{g}\perp\Delta$triangleright $\mathbf{a}\perp\mathbf{b}$ denotes orthogonal projection of $\mathbf{a}$ onto $\mathbf{b}$
17:     end if
18:     $\mathbf{\theta}\leftarrow\mathbf{\theta}_{\text{ref}}$. $\triangleright$ Restore original state.
19:     $\theta\leftarrow\mathbf{O}(\theta,\mathbf{g})$triangleright Take step with main optimizer.
20:end for
At a high level, an implementation of PROFIT involves defining an optimizer wrapper which takes as argument two separate standard optimizers, $\mathbf{O}$ and $\mathbf{O}^{(\text{ref})}$ as inputs. The latter “reference” optimizer perturbs the system from equilibrium, while the former “main” optimizer uses this perturbation to make a final update. The entire logic of PROFIT can be encapsulated within the logic of this optimizer wrapper class, making the method very portable, modular, and invariant to the choice of “main” optimizer.
Given a model $\mathcal{M}$ with weights $\theta$ which has been trained on data $\mathbf{X}$ corresponding to an old task, and a data source $\mathbf{X^{\prime}}$ from a similar domain corresponding to a new task (with a task loss $L$), we aim to fine-tune $\mathcal{M}$ to work well on both the old and new tasks. Note that we make a strict assumption about having a converged model $M$ available as an input to PROFIT; any random or untrained weights in the model can lead to poor performance. Additionally, our model also requires two instantiated optimizers, a standard optimizer $\mathbf{O}$ and a reference optimizer $\mathbf{O}^{(\text{ref})}$. The user can tune $\mathbf{O}$ as per their needs, while we recommend using SGD for $\mathbf{O}^{(\text{ref})}$.
A step through our optimizer works as follows. First, we store the current state of $\mathcal{M}$ by saving $\theta$ into $\theta_{\text{ref}}$. Next, we draw $n_{\text{ref}}$ batches from $\mathbf{X^{\prime}}$, one at a time, and iteratively minimize $L$ with the reference optimizer $\mathbf{O^{(\text{ref})}}$. We have now perturbed the system from equilibrium and must decide the best way to restore the said equilibrium.
To do so, we first calculate $\Delta:=\theta-\theta_{\text{ref}}$, the total displacement of our original position after $n_{\text{ref}}$ steps of the optimizer $\mathbf{O^{(\text{ref})}}$. We reason that if the stored equilibrium state corresponds to a good critical point of the original model, then $\Delta$ corresponds to the benign gradient direction that will restore the original critical point. We thus have two potentially conflicting gradient updates: the one corresponding to $\Delta$, and the one corresponding to the next queried update by $\mathbf{O^{(\text{ref})}}$, which we call $\mathbf{g}$. We need to decide on how to take a single gradient step that is consistent with both options.
We then borrow an idea from PCGrad [55], which reconciled conflicting gradients by orthogonally projecting them onto each other in a pairwise fashion. Crucially, we choose to only project $\mathbf{g}$ onto $\Delta$, and not the other way around, because $\Delta$ represents a gradient on the old dataset which may no longer be accessible and, therefore, must be treated with more care. Therefore, we end with the two gradient updates $\mathbf{g}$ and $\mathbf{g}\perp\Delta$, and we take both steps by first restoring $\theta\mapsto\theta_{\text{ref}}$ and then allowing $\mathbf{O}$ to take a step in the $\mathbf{g}\perp\Delta$ direction. This process is repeated until training is complete.
Interestingly, this formulation allows us to view fine-tuning as temporal multi-task learning, with the two tasks being: (1) pretraining ($\Delta$), and (2) fine-tuning ($\mathbf{g}$). To the best of our knowledge, this is the first time it has been viewed through this lens, and this insight may pave avenues for future research on fine-tuning.
3.2 Hyperparameter Discussion and Tradeoffs
PROFIT introduces three main hyperparameters for fine-tuning: $n_{\text{ref}}$, $\mathbf{O}^{\text{(ref)}}$, and $\lambda_{\text{ref}}$. As previously discussed, $n_{\text{ref}}$ controls the degree of exploration for the reference optimizer, while $\lambda_{\text{ref}}$ controls the step size at each iteration of the reference step. The choice of $\mathbf{O}$ can be set by the practitioner based on their needs, but $\mathbf{O}^{(ref)}$ should generally be set to standard SGD, as the gradient calculations that drive PROFIT are cleanest when the reference updates are simple.
The primary concern for practitioners using PROFIT is the cost of setting $n_{\text{ref}}$, as it requires $n_{\text{ref}}$ additional optimization steps per training step. In practice, the practitioner is encouraged to start with $n_{\text{ref}}=1$, which will often work well, and only to increase the value if they observe the performance staying flat. However, we note two reasons why the potentially increased compute of PROFIT is not a large issue: (1) we find that PROFIT generally converges quicker, possibly due to the positive regularization effects of the method, and (2) fine-tuning is generally run for much fewer steps than training from scratch, so the effects of this additional training time are often still very tractable. For all experiments, we use $n_{\text{ref}}=1$.
Mathematically, $\lambda_{\text{ref}}$ encodes how quickly we travel from equilibrium to study the general shape of the loss surface, so that we can make an informed decision on where to go next. In practice, we find that there is often a sweet spot somewhere between $\lambda_{main}/10$ and $\lambda_{main}/10000$, but the exact value is heavily dependent on the loss surface for the specific problem. It may in fact turn out that the optimal value of $\lambda_{\text{ref}}$ tells us something about the fundamental properties of that particular setting’s loss surface, but such analysis falls outside the scope of the present work.
3.3 Theoretical Considerations
We now list a few simple theoretical properties of PROFIT, along with sketches of proofs. We first prove that PROFIT is “correct”, in the sense that it is able to decrease the loss on the old model task/setting. We also add a discussion of when PROFIT might fail by enumerating all stable points of PROFIT, but argue why this is generally not an issue in practical settings.
{theorem}
[] (Correctness on old data.) Take a model $\mathcal{M}(\mathbf{x};\theta)$ converged on data $\mathcal{X}_{\text{old}}$ with loss $L_{\text{old}}$, which we would now like to fine-tune on data $\mathcal{X}_{\text{new}}$ with loss $L_{\text{new}}$. Suppose that $\mathbf{O}^{(\text{ref})}$ takes $\theta\mapsto\theta^{\prime}$. A single step of PROFIT on batch $\mathbf{x}_{\text{new}}$ with a sufficiently small learning rates $\lambda_{\text{main}}$, $\lambda_{\text{ref}}$ will decrease $L_{\text{old}}(\mathcal{X}_{\text{old}})$ from its value at $\theta^{\prime}$.
Proof Sketch: If $\mathcal{M}$ is converged on data $\mathcal{X}_{\text{old}}$, then it must be at a local minimum of the old data. If the learning rate is sufficiently small, then $\nabla_{\theta^{\prime}}L_{\text{old}}$ must point in the $\Delta$ (as defined in Alg 1) direction, and so an update with the gradient $\Delta+\langle\Delta,\mathbf{g}\rangle$ will necessarily decrease $L_{\text{old}}$.∎
The prior theorem establishes that PROFIT accomplishes precisely what it seeks to do: even if the old data is no longer available, PROFIT allows us to train as if we can still compute the full loss function of the old data. As the system moves further away from the old equilibrium, we are able to restore some of the function of that equilibrium through these regularized updates. Notably, there is a case where PROFIT leads to a trivial update.
{theorem}
[] (Stable points.) Suppose a model has weights $\theta$ and $\mathbf{O}^{(\text{ref})}$ maps $\theta\mapsto\theta^{\prime}$ and $\mathbf{O}^{(ref)}$ is SGD. If we are not at a critical point of $L_{\text{new}}$, PROFIT will result in zero change in the model weights $\mathbf{w}$ if and only if ^$\hat{\nabla}_{\theta}L_{\text{new}}=\hat{\nabla}_{\theta^{\prime}}L_{\text{new}}$, where $\hat{\nabla}$ refers to the unit vector corresponding to $\nabla$.
Proof: $\Delta$ in this case would point in the direction $-\nabla_{\theta}L_{\text{new}}$, at which point $\nabla_{\theta}L_{\text{new}}\perp\Delta=0$, and the total update from PROFIT will be just a restoration to $\theta$. If the gradient equality does not hold, $\nabla_{\theta}L_{\text{new}}\perp\Delta\neq 0$ will lead to a non-zero total update. ∎
{corollary}
[] (Linearity forces a stable point.) In the situation defined by Theorem 3.3, PROFIT will encounter a stable point if the loss surface is perfectly linear between $\theta$ and $\theta^{\prime}$. ∎
The prior theorem and corollary demonstrate that PROFIT will fail to move the system when the loss surface becomes exactly locally linear at a point. This almost never occurs for high-dimensional loss surfaces that exist for deep models, and even in the 2-dimensional toy example described in Section 4.1 we see a reasonable performance.
3.4 Assumptions and Limitations
The key assumptions made by PROFIT are:
  1. 1.
    Proximal Fine-tuning: pre-training and fine-tuning datasets come from similar distributions or modalities.
  2. 2.
    Well-trained: the initial network is converged and not randomly initialized.
Orthogonalizing between $\mathbf{g}$ and $\Delta$ implies that we are interested in resolving the conflict between the original and fine-tuning datasets. If the original and fine-tuning distributions are significantly different, the orthogonalized gradient direction may not benefit either tasks and lead to destructive interference. Therefore, our interpretation holds only if the original and fine-tuning datasets are from proximal distributions. Assumption 1 may differ from typical fine-tuning literature and appear greatly limiting, but is exceedingly common in practice. For instance, a self-driving car may need to train a classification head for a new type of scooter, while also maintaining good performance on previously observed vehicles.
Assumption 2 is required since our method is exclusively designed for fine-tuning (which is a guiding principle). Nonetheless, we include it here to emphasize to the reader that PROFIT cannot be used for general model optimization.
As discussed in Sec 3.3, PROFIT requires inference on $n_{\text{ref}}+1$ batches (two for $n_{\text{ref}}=1$) per iteration, hence is slower than standard optimizers for fine-tuning. In practice, we observe that PROFIT converges earlier than corresponding baselines (which are allowed to train longer for fair comparison), which may mitigate this. We also note that typical fine-tuning settings are shorter and not as involved as training from scratch, which may further alleviate this concern. Another limitation is that our method would consume slightly more memory as a result of instantiating two optimizers ($\mathbf{O}^{\text{(ref)}}$ and $\mathbf{O}$).
4 Experiments
We now detail a number of experiments for PROFIT in diverse settings: low-dimensional toy setting to image classification, visual task adaptation, and large-scale motion prediction. We primarily focus on comparisons to standard fine-tuning with standard optimizers (on either the full model or just the model head), as those are - by a large margin - still the most commonly used fine-tuning methods in the industry due to their known performance and ease of implementation. We will show that PROFIT, while being easy to implement, provides a significant performance boost in all cases.
4.1 A Simple Toy Example
We present results in this section on a simple, low-dimensional toy example. We find that testing new algorithms in low-dimensional settings is always instructive and provides us with insight that cannot be gleaned from the noise of higher-dimensional experiments.
Namely, we fit a number of simple MLP models onto a 2D function. We pick the 2D function $f(\mathbf{x})=\sin(10|\mathbf{x}|)$, because both radially symmetric and periodic functions are generally more challenging for neural networks to fit. We further add $\mathcal{N}(0,1)$ noise to the output. Even though we want to fit to a low-dimensional example, it is still important for the problem to be difficult enough to see interesting behavior within our models, especially given the dimensionality requirements as discussed in Section 3.3. The original dataset consists of input-output pairs where input coordinates are drawn independently from $\mathcal{U}[-1.0,1.0]$, while the new dataset we wish to fine-tune on has input coordinates drawn from $\mathcal{U}[0.8,1.5]$. This is clearly also challenging because, although the domains of the two datasets overlap in the interval $[0.8,1.0]$, they are largely nonoverlapping.
Our MLP consists of three layers with weights of shape $[2,500]\rightarrow[500,500]\rightarrow[500,1]$, and we use RMSProp as our baseline optimizer. The baseline model is trained on the original data split only for 10000 steps at a learning rate of 1e-2, with fine-tuning runs trained at a learning rate of 5e-4 for 1500 steps. For exact details on the training procedure, please refer to the supplementary material.
Figure 2: Fine-tuning toy example visualizations. The top row shows ground truth distributions for the original, new, and combined datasets, while the bottom row depicts model predictions for different training strategies: training on original data only, head-only fine-tuning, full model fine-tuning, and using the proposed PROFIT optimizer for full model fine-tuning. This highlights the PROFIT optimizer’s effectiveness in retaining old task knowledge while adapting to new tasks. Best viewed in color.
Method Original Data Error ()($\downarrow$)( ↓ ) New Data Error ()($\downarrow$)( ↓ )
Baseline Trained on Original Data 0.0054 1.907
fine-tune on New Data (Full Model) 0.705 0.504
fine-tune on New Data (Head Only) 0.110 0.572
PROFIT on New Data 0.046 0.501
Table 1: Results on a 2D fitting problem. A baseline is trained on just the original data domain and then fine-tuned at a lower learning rate on a new data domain (both the full model and the head only). This is compared to PROFIT, which is trained on the full model. As expected, there is a clear tradeoff between training on the full model versus the head only; the latter provides some protection against performance regression on the original data, but also is less able to adapt to the new data. PROFIT outperforms both baselines and shows impressive resilience in maintaining performance on the original split. All results have standard error $\leq 0.01$.
The results are visualized in Figure 2. Visually, the benefits of PROFIT are fairly pronounced; fine-tuning on the full model creates a warped shape bearing minimal resemblance to the original ground truth, and even fine-tuning on the head layer only still produces sizable shifts in the overall height of the output shape. PROFIT effectively remembers the original shape with only minimal regressions along the steep edges of the distribution.
Numerically, as shown in Table 1, after fine-tuning with different techniques, it is clear that PROFIT outperforms the baseline in not forgetting the original dataset distribution. Even though PROFIT modifies all model weights, it mitigates forgetting relative to the head-only fine-tuning baseline by a sizable margin, even though PROFIT acts on significantly more weights. Fine-tuning of the full model leads to disastrous results, with significant deformations of the predictions. PROFIT allows us the flexibility of full-model fine-tuning without the drawbacks of catastrophic forgetting.
Even though PROFIT is partially inspired by LWF [30], it is worth noting that LWF is ill-suited for shifting dataset domains, as it uses evaluation on the new data with the old model to generate additional supervision. With large input domains shifts, this supervision would be of very poor quality. Thus, these experiments also serve to show that PROFIT is a quite general fine-tuning tool.
Method Optimizer CIFAR10 Acc ($\uparrow$) CIFAR100 Acc ($\uparrow$)
ResNet-18 SGD 24.49% 74.59%
Adam [23] 34.64% 73.11%
Lookahead [58] 26.10% 73.82%
Adan [54] 35.17% 72.70%
PROFIT 35.26% 74.70%
ViT-Tiny SGD 53.63% 61.98%
Adam [23] 56.00% 58.99%
Lookahead [58] 55.64% 61.35%
Adan [54] 53.04% 61.62%
PROFIT 58.53% 62.20%
ViT-Small SGD 57.93% 65.04%
Adam [23] 58.60% 63.93%
Lookahead [58] 57.81% 65.27%
Adan [54] 53.85% 64.09%
PROFIT 59.02% 65.44%
Table 2: Results on image classification benchmarks after fine-tuning on CIFAR100. There is notable improvement for PROFIT on both CIFAR10 and CIFAR100 benchmarks, indicating that our method is not only able to perform well on the fine-tuning task, but also alleviates forgetting on the pretraining task.
Method $n_{\text{ref}}$ CIFAR10 Acc ($\uparrow$) CIFAR100 Acc ($\uparrow$)
ResNet-18 1 35.26% 74.70%
2 32.55% 73.57%
5 39.27% 71.42%
ViT-Tiny 1 56.75% 62.35%
2 53.10% 61.67%
5 56.56% 59.14%
ViT-Small 1 59.02% 65.44%
2 58.58% 64.91%
5 56.54% 63.75%
Table 3: Ablation study on $n_{\text{ref}}$ for image classification for different backbones.The results demonstrate that varying $n_{\text{ref}}$ can have a significant impact on model accuracy when using PROFIT.
Method $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ CIFAR10 Acc ($\uparrow$) CIFAR100 Acc ($\uparrow$)
ResNet-18 10 35.26% 74.70%
100 35.16% 74.45%
1000 34.35% 74.27%
10000 35.83% 74.41%
ViT-Tiny 10 57.32% 61.99%
100 56.75% 62.35%
1000 58.53% 62.20%
10000 53.78% 62.05%
ViT-Small 10 58.41% 66.29%
100 59.02% 65.44%
1000 56.96% 65.53%
10000 56.84% 65.95%
Table 4: Ablation study on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ for PROFIT in image classification for different choices of backbone architectures. For ResNet-18, smaller values of $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ yield relatively stable accuracy, while ViT-based models show a more pronounced change, with peak performance observed at specific ratios. This result highlights the importance of fine-tuning $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ to achieve optimal performance for different backbone architectures in image classification tasks.
4.2 Image Classification
Natural Specialized Structured
Cifar100
Caltech101
DTD
Flower102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Ele
Full [21] 68.9 87.7 64.3 87.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1
Linear [21] 64.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.5 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
PROFIT 58.9 55.6 51.4 92.4 50.8 19.6 37.7 79.3 53.8 43.5 73.5 12.6 37.9 21.1 77.5 10.2 24.6 23.6 10.7
PROFIT (warmup) 69.4 90.5 69.6 98.9 89.2 89.4 52.0 84.9 95.7 85.6 74.3 80.5 64.7 51.9 80.6 82.3 49.3 32.7 35.1
Table 5: Image classification accuracy on each task in the VTAB-1k benchmark. We compare PROFIT against standard fine-tuning strategies, including full fine-tuning and linear fine-tuning (denoted by Full [18] and Linear [18]). Results are reported across various datasets categorized into Natural, Specialized, and Structured domains. A naive implementation of PROFIT performs suboptimally due to the violation of the proximity condition in this setting, but when supplemented with AdamW warm-up PROFIT (warm-up) demonstrates superior performance and consistently achieves higher accuracy scores in multiple datasets. For each block (Natural, Specialized, and Structured), the highest accuracy are highlighted in bold.
Next, we demonstrate the effectiveness of PROFIT on the image classification task. CIFAR10 and CIFAR100 [28] are the de-facto benchmarks for image classification. The dataset consists of $32\times 32$ input images, and the task for the network is to classify an image into one of $K$ categories.
We first train a network on CIFAR10 (pre-training), and fine-tune the network on CIFAR100. As discussed earlier, one of the assumptions for PROFIT is that of proximality, where both the pretraining and fine-tuning datasets need to be from a similar domain distribution. This assumption is met for our choice of CIFAR10 and CIFAR100, as they are both labeled subsets of Tiny Images [49].
We experiment with various choices of backbones, including ResNet-18 [17], ViT-Tiny, and ViT-Small [12]. We compare PROFIT against a variety of the baselines, such as Stochastic Gradient Descent (SGD), Adam [23], Lookahead [58], and Adan [54]. We run the experiments 3 times and report the best results, but see low variance in our results.
Our results are shown in Table 2, which show that PROFIT is able to perform better on the fine-tuning task across all backbone networks and optimizers. We also show better CIFAR10 accuracy after fine-tuning, indicating that our method is better at maintaining old task performance as well. For example, fine-tuning a ResNet-18 with Adan [54] optimizer leads to 35.17% accuracy on CIFAR10 and 72.70% accuracy on CIFAR100, while using PROFIT shows better original task performance (35.26% accuracy) and fine-tuning task performance (74.70% accuracy). With ViT-Tiny, Adam fine-tuning results in 55.64% accuracy on CIFAR10 and 61.35% accuracy on CIFAR100, but our method comfortable outperforms with 58.53% on CIFAR10 and 62.20% on CIFAR100. fine-tuning ViT-Small using Adam shows 58.60% accuracy on CIFAR10 and 63.93% accuracy on CIFAR100, however, PROFIT shows better performance with 59.02% accuracy on CIFAR10 and 65.44% accuracy on CIFAR100.
We ablate $n_{\text{ref}}$ in Table 4. In general, we observe that increasing $n_{\text{ref}}$ leads to better performance on the original task while compromising performance on the fine-tuning task. This follows from our discussion in Sec 3.2, since $n_{\text{ref}}$ controls the degree of exploration away from the stable point. Nonetheless, we recommend $n_{\text{ref}}=1$ as a starting point.
We also ablate performance on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ in Table 4, which shows the best choice may vary for the choice of backbone. In general, larger values of $\lambda_{\text{ref}}$ promotes new task accuracy, while smaller values mitigate catastrophic forgetting effectively. These results are reasonable, as smaller values of $\lambda_{\text{ref}}$ equates to the reference point lying closer to the original model state. However, we always were able to beat the baseline on both old and new task accuracies simultaneously.
We also use this setting to substantiate Assumption 2 (Sec 3.4), which states PROFIT does not work well without a converged model. When we train a ResNet-18 from scratch on CIFAR100 using PROFIT, we get 1.05% accuracy, which is as good as random guessing.
4.3 Visual Task Adaptation Benchmark
The Visual Task Adapation Benchmark (VTAB-1K) [56] is a popular representation learning benchmark to evaluate generalization across a diverse set of classification tasks. It consists of 19 different tasks from a broad spectrum of domains, with the aim to learn feature representations which perform well on all of these tasks. Performance is measured by computing classification accuracy on each of these 19 tasks.
We use ViT-Base [20] with ImageNet pretraining as the backbone for our experiments. This setting is an example of a setting where the original (ImageNet) and fine-tuning distributions (VTAB-1K) significantly differ, violating Assumption 1 (Sec 3.4). While our method is not well tuned for these fine-tuning settings, this situation is common in practice. In this section, we show how a practitioner can leverage our method in such settings.
We show our results in Table 5, and compare our method with AdamW [34] fine-tuning of the entire model (Full – row 1) as well as final layer fine-tuning (Linear – row 2). Full fine-tuning using PROFIT does not work well, showing poor performance across all 19 tasks in the benchmark.
However, we provide a training recipe on how PROFIT can be used in such settings. First, we fine-tune the model towards the target distribution with an AdamW warmup for 10 epochs, and then apply PROFIT to fine-tune for the remaining 90 more epochs (PROFIT(warmup)). Intuitively, this moves the model’s distribution somewhere between the pre-training and fine-tuning distribution, which makes it amenable to the structure of PROFIT. This approach outperforms full model fine-tuning with AdamW. So while our method as expected is not recommended for non-proximal settings (as discussed in Section 3.4), using a short warmup phase with another optimizer allows PROFIT to substantially outperform full fine-tuning with standard optimizers.
4.4 Large-Scale Robotics Motion Prediction
We now show how PROFIT performs on an extremely challenging benchmark with a large, high-dimensional dataset. The Waymo Open Motion Dataset (WOMD)  [13] is a large-scale driving dataset collected from realistic scenarios. The task is to predict future trajectories of an agent over the next 8 seconds, given multi-modal observations in the last second, including the agent’s history, nearby agent histories, map information, and traffic light states. The prediction model follows a state-of-the-art early-fusion transformer, Wayformer [38], by fusing multi-modal input features through a self-attention transformer and predicting future trajectory samples using learned latent queries.
We first train a model on car trajectory prediction, and then fine-tune the model on data from the same class (car) as well as different classes (pedestrian). For our baselines, we use the AdamW [34] optimizer. WOMD offers an important difference from previous settings in that we can see how well PROFIT performs on (1) fine-tuning a model on the same data, and (2) fine-tuning a model on a different domain-shifted task. For instance, CIFAR100 is explicitly part of the same distribution as CIFAR10, but pedestrian trajectories are semantically different than car trajectories.
Figure 3: Results for PROFIT on Waymo Open Motion Dataset. (a) Error curves through training tabulated for FDE at 3s and 8s (top) and ADE at 3s and 8s (bottom). PROFIT outperforms both fine-tuning baselines by a sizable margin. (b) Visualizations of motion prediction outputs for both the baseline fine-tune model (top) and PROFIT (bottom). Trajectory ground truth is shown as a shaded bar and denser lines represent more confident predictions. Although the differences are often subtle, PROFIT generally produces more confidently correct predictions. Best viewed in color.
Method Target Class ADE@3s (m) ADE@8s (m) FDE@3s FDE@8s
Baseline - 0.461 1.327 1.024 2.581
fine-tune (F) Car 0.458 1.322 1.021 2.548
fine-tune (H) Car 0.456 1.303 1.009 2.507
PROFIT Car 0.454 1.299 1.008 2.489
fine-tune (F) Ped 0.214 0.621 0.465 1.242
fine-tune (H) Ped 0.232 0.724 0.508 1.544
PROFIT Ped 0.203 0.579 0.427 1.145
Table 6: Results on Waymo Open Motion Dataset. F stands for full-model fine-tuning and H stands for head (last layer) only. Standard errors are within 0.005m for 3s metrics and 0.015m for 8s metrics. There is a minor but noticeable improvement for PROFIT on the car-to-car benchmarks and a substantial improvement for PROFIT on the car-to-ped benchmarks.
Figure 4: We compare the baseline with PROFIT on an example from DriveLM. The model fine-tuned with PROFIT is able to perceive the traffic light and black sedan in the scene, while the baseline (fine-tuned with AdamW) does not detect the traffic light and hallucinates the presence of a white truck. Consequently, the baseline suggests running the red light, while our method follows traffic rules by staying stationary. Best view in color.
The results are shown in Figure 3 and Table 6. Average Distance Error (ADE) measures the average distance between ground truth and prediction at each point of the predicted trajectories, while Final Distance Error (FDE) only considers the error in the final point of the predicted trajectories. We see that in all cases, PROFIT outperforms the baseline, with sizable benefits in the car-to-pedestrian fine-tuning task. Consistent with the results in other settings, fine-tuning with head only provides subpar performance on the car-to-pedestrian task, while full fine-tuning does reasonably well but is still outstripped by PROFIT. We conclude that PROFIT leads to a better results to re-purpose a model for a new, correlated setting.
Backbone Optimizer Accuracy (%) $\uparrow$ GPT Score (%) $\uparrow$ Match Score (%) $\uparrow$ BLEU-4 $\uparrow$ ROUGE_L $\uparrow$ CIDEr $\uparrow$ Final Score $\uparrow$
LLAMA-Adapter-v2 [15] AdamW 62.21 70.57 36.33 0.541 7.16 0.023 56.98
PROFIT 67.88 72.12 37.37 0.557 7.28 0.035 59.16
Table 7: Results on the DriveLM Benchmark. We fine-tune LLAMA-Adapter-v2 [15] and compare PROFIT with AdamW, the de-facto method for fine-tuning VLMs. Our method shows noticeable improvements on all metrics, showcasing the applicability of PROFIT to large foundational models.
Backbone $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ Accuracy (%) $\uparrow$ GPT Score (%) $\uparrow$ Match Score (%) $\uparrow$ BLEU-4 $\uparrow$ ROUGE_L $\uparrow$ CIDEr $\uparrow$ Final Score $\uparrow$
LLAMA-Adapter-v2 [15] 10 65.82 71.99 37.21 0.550 7.21 0.029 58.54
100 65.83 72.19 37.16 0.550 7.22 0.029 58.64
1000 65.74 72.07 37.02 0.550 7.22 0.030 58.54
10000 67.88 72.12 37.37 0.557 7.28 0.035 59.16
Table 8: Ablation study on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ for PROFIT on the DriveLM Benchmark. Higher values of $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ improves perception and prediction accuracy and language scores, showcasing the strength of PROFIT in generalizing to both VQA and scene understanding tasks.
We also see minor but noticeable improvements in the car-to-car benchmark, suggesting that PROFIT can be used to extract more performance from any model that has already converged. We could even imagine a scenario in which regular additional training runs using PROFIT are enabled as a standard model maintenance practice.
4.5 Driving with Vision Language Models
DriveLM [46] is a visual question-answering (VQA) benchmark to evaluate vision-language models (VLMs) for an autonomous driving setting. The DriveLM-nuscenes dataset contains sensor inputs and a text prompt, and the network aims to provide accurate textual responses for prompts which probe perception, prediction, and planning capabilities respectively.
The DriveLM benchmark leverages a variety of VQA metrics for evaluation. BLEU [39] evaluates precision by measuring n-gram similarity between generate and reference texts. ROUGE_L [31] evaluates recall by measuring the longest common subsequence between the generated and reference texts. CIDEr [51] evaluates quality by computing the cosine similarity between the n-gram TF-IDF vectors for the generated and reference sentences. GPT Score aims to captures semantic nuances missed by aforementioned metrics by using ChatGPT (GPT-3.5-Turbo) as an evaluator. In addition to these metrics, perception accuracy is evaluated using the ground-truth objects in the scene, while prediction accuracy is evaluated over discretized future states. Match score evaluates whether the VLM correctly understands the order in which to attend to other agents in the scene. The final score for this benchmark is a weighted average of these metrics. For further details, we refer the reader to the evaluation criteria for the DriveLM Challenge [11].
We use a pre-trained LLAMA-Adapter-v2 [15], which performs bias tuning on LLAMA-7B [50]. We compare PROFIT with AdamW [34], the de-facto method for fine-tuning VLMs.
We report our results on the validation set of DriveLM in Table 7. PROFIT improves perception and prediction accuracy over AdamW by 5.6%, GPT score by 1.5%, Match Score by 1%, and the Final Score by 2. In addition, PROFIT provides better outputs for VQA, as seen by improved performance on language metrics (BLEU-4, ROUGE_L, CIDEr). We provide an inference visualization of our method in Fig 4. The baseline, fine-tuned with AdamW, is unable to detect the traffic light and black sedan in the scene. This leads to the suggested plan of running the red light, which is a potentially catastrophic error. On the other hand, PROFIT successfully identifies the red light and follows traffic regulations by planning to remain stationary. We conclude PROFIT is a useful tool in the era of VLMs to extract better performance.
We ablate performance on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$, a key hyper-parameter to PROFIT, in Table 8. In general, larger values encourage exploration and improves fine-tuning task accuracy. This follows Sec 4.2, since smaller ${\lambda_{\text{ref}}}$ implies the reference step away from the pre-trained model is smaller.
5 Conclusion
We propose PROFIT, an optimizer which is designed to robustify models during fine-tuning. Unlike standard optimizers which cannot assume much about a training setting, our ability to assume confidence about the prior state of a model allows PROFIT to act as an effective regularizer and prevent a model’s weights from diverging too much from its previously good state. We showed that PROFIT performs well in various settings: (1) fine-tuning to new data off the data manifold in our toy example, (2) fine-tuning to new CIFAR100 data from CIFAR10 pretraining, (3) fine-tuning to a new data distribution in VTAB-1K, and (4) fine-tuning a model on a completely new task as well as on the same exact task on large-scale motion prediction. In all cases, PROFIT surpassed the performance of the near-ubiquitously used, standard fine-tuning approaches, while being just as easy to integrate within a generic model training pipeline. We believe that PROFIT should not only be an important tool in the deep learning practitioner’s arsenal but opens the discussion around developing a new class of optimizers that will crucially support the new age of deep learning where fine-tuning becomes the primary training paradigm.
6 Acknowledgment
We thank Carl Vondrick, David Hayden, Mao Ye, Navaneeth Bodla, Amirreza Shaban, Greg Meyer, and Sidd Srinivasa for their guidance and discussions which help shape our work.